OpenStudy (e.cociuba):

For each quadratic equation, find the: a.) Roots b.) y-intercept c.) Vertex y= 4(x-4)^2 +16

OpenStudy (anonymous):

c?

OpenStudy (e.cociuba):

Ok so for the root part I can put it in my calc and find the roots. Give me a sec...

OpenStudy (campbell_st):

roots occur when y = 0, so let y = 0 and solev for x so you have 16 = 4(x -4)^2 divide by 4 4 = (x -4)^2 take the square root of both sides \[\pm 2 = x - 4\] add 4 to both sides \[x = -2 + 4....and.....x = 2 + 4\] just simplify for your roots. y - intercept is when x = 0 and your equation is in vertex form \[y = (x - h)^2 + k\] (h, k) is the vertex.... just match them up.

OpenStudy (e.cociuba):

Ok thanks that makes so much more sense!

OpenStudy (e.cociuba):

Ok so for the roots it's gunna be (2,6) Correct?

OpenStudy (e.cociuba):

SO the y-intercept is 80 correct?

OpenStudy (campbell_st):

the roots are correct. The y-intercept x = 0 \[y = ( 0 - 4)^2 + 16 = 16 + 16 \] y intercept is y = 32

OpenStudy (e.cociuba):

y= 4(0-4)^2 +16: I put that in my calculator and got 80. I think u forgot to put the 4 in front of the parenthesis....